OpenStudy (lasttccasey):

All edges of a cube are expanding at a rate of 5 centimeters per second. (a) How fast is the volume changing when each edge is 1 centimeter(s)?

OpenStudy (lasttccasey):

\[V=s^3\] I can get derivative and plug in but I don't get the correct answer?

OpenStudy (lasttccasey):

\[\frac{dV}{dt}=2s^2\frac{ds}{dt}\] should be the derivative but I don't know why I'm getting it wrong

OpenStudy (cwrw238):

dV/dt = 3s^2 * ds/dt

OpenStudy (cwrw238):

= 3 * 5

OpenStudy (lasttccasey):

cant believe I put down a 2 -_- thanks

OpenStudy (cwrw238):

yw